OpenStudy (anonymous):

PLEASE HELP! find the equation of the line in point slope form that is parallel to y+1=3(x-1) and passes through and passes through (2,6)

OpenStudy (mrnood):

first expand the bracket then rewrite the equation so it reads y=???? This will give you more information about the line.... Post your result here - I'll help more

OpenStudy (anonymous):

y=mx+b?

OpenStudy (mrnood):

that's what you're tryiong to get to - but you need the real numbers. do the 2 steps I wrote above....

OpenStudy (anonymous):

so i would plug in 2,6 in the equation?

OpenStudy (mrnood):

no @M4thM1nd - PLEASE do not post the whole answer

OpenStudy (anonymous):

ok

OpenStudy (mrnood):

I have given you 2 steps above - PLEASE do them - they will get you on the way to an answer first expand the bracket then rewrite the equation so it reads y=????

OpenStudy (anonymous):

both lines must be parallel... so they must have the same "m"

OpenStudy (anonymous):

after that, you have a point (2,6) and the "m", now just write the equation of line

OpenStudy (anonymous):

y=3x?

OpenStudy (anonymous):

Yes!

OpenStudy (mrnood):

Why don't you PLEASE do exactly the 2 steps I posted? IF I didn't think they would help you to get the answer then I wouldn't have suggested them

OpenStudy (anonymous):

what do you mean by expand the bracket

OpenStudy (mrnood):

3(x+1) = ??

OpenStudy (anonymous):

so then i would plug in 2,6?

OpenStudy (mrnood):

I give up

OpenStudy (anonymous):

lol im reeally bad at this

OpenStudy (mrnood):

first expand the bracket means 3(x+1) = then rewrite the equation so it reads y=????

OpenStudy (anonymous):

y=3(x+1)

OpenStudy (mrnood):

y+1= 3(x+1) is the equation given. NOW write it as y+1= ???? where ??? = the expanded bracket (i.e. multiply the 3 into all terms inside the bracket)

OpenStudy (anonymous):

y+1=3x+3?

OpenStudy (mrnood):

good! NOW rearrange that equation so it is y=???

OpenStudy (anonymous):

y=3x+3?

OpenStudy (mrnood):

what happened to the 1 from y+1?

OpenStudy (anonymous):

so 1=3x+3?

OpenStudy (mrnood):

I'm sorry to be blunt - but if you cannot manipulate equations accurately then you cannot do this type of question. Please revise re-arranging equations

OpenStudy (anonymous):

i really dont get it thanks for the help i give up

OpenStudy (anonymous):

don't give up, let me try to make you understand

OpenStudy (anonymous):

you have this equation y+1 = 3*(x-1) , right?

OpenStudy (anonymous):

what is the slope of this line?

OpenStudy (anonymous):

i skipped the question already can we try a new one?

OpenStudy (anonymous):

let's try this one first, because if you don't understant this one, probabily you won't solve the next one

OpenStudy (anonymous):

okay so the slope is 3?

OpenStudy (anonymous):

can you identify what is the slope of this equation: y+1=3*(x-1)

OpenStudy (anonymous):

that's right

OpenStudy (anonymous):

whats next?

OpenStudy (anonymous):

there is many forms of the equation of a line, let's use this one: (y - y0) = m*(x - x0) where the point (x0, y0) is a point IN this line (can be any point in this line), and m is the slope of this line

OpenStudy (anonymous):

till here you understantd?

OpenStudy (anonymous):

okay so where the m is ill put 3?

OpenStudy (anonymous):

wait, this is a general form, the Problem asks you to find the equation of the line that is PARALLEL to (y+1)=3*(x-1) and passes through point (2,6)

OpenStudy (anonymous):

if two lines are parallel they must have the same slope right? so the same m

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok, so now we know that the line that we want to find has slope 3, i.e m=3

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

now we use the general equation of a line: (y - y0) = m*(x - x0) Now m = 3 and one point that we want that this line passes through is (2,6), so x0 = 2 and y0 = 6

OpenStudy (anonymous):

now just plug this values, and we get (y - 2) = 3*(x - 6)

OpenStudy (anonymous):

and that is the answer

OpenStudy (anonymous):

sorry (y - 6) = 3*(x - 2)

OpenStudy (anonymous):

i got it now well no expert yet

OpenStudy (triciaal):

parallel lines have the same slope. the slope of the line is 3 now you know it passes through (2, 6) so when x = 2 y = 6 m= 3 you can find b y = mx + b is the equation of the line in slope-intercept form 6 = 3(2) + b so b = 0 for any point on the line (x, y) y = 3x + 0 so the line parallel is y = 3x

OpenStudy (anonymous):

Good luck with your study

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

@M4thM1nd would it be different for slope intercept form